A bird watcher meanders through the woods, walking 0.40 km due east, 0.70 km due south, and 2.15 km in a direction 31.0° north of west. The time required for this trip is 3.00 h.

(a) Determine the magnitude and direction (relative to due west) of the bird watcher's displacement. Use kilometers and hours for distance and time, respectively.

Magnitude: ___km

Direction:_____north of west



(b) Determine the magnitude and direction (relative to due west) of the bird watcher's average velocity.

Magnitude: _____km

Direction:_____ north of west
Answer #1

net displacement along x axis

dx = 0.4 - 2.15 cos 31 = - 1.44 km

net displacement along y axis

dy = - 0.7 + 2.15 sin 31 = 0.4073 km

net displacement

d^2 = dx^2 + dy^2

d^2 = 1.44^2 + 0.4073^2

d = 1.5 km

direction,

x = arctan ( 0.4073 / 1.44) = 15.793 north of west.

=======

b)

av velocity = d/t = 0.5 km/h

direction same as (a)

=====

Comment before rate in case any doubt, will reply for sure.. goodluck